Rice Company has a unit selling price of $690, variable costs per unit of $390, and fixed costs of $286,300. Compute the break-even point in units using (a) the mathematical equation and (b) unit contribution margin. (Round answers to 0 decimal places, e.g. 1,225.)

Whats the break even point a)mathematical equation: _________ units B) Unit contribution margin __________ units
Answer #1

A) At Breakeven sales

Contribution margin - Fixed cost = 0

(690-390)*S - 286,300 = 0

300S = 286,300

Breakeven point = 286,300/300

= 954 units

B) Unit contribution margin = 690-390 = 300

Breakeven units = Fixed cost /Unit contribution margin

= 286,300/300

= 954 units